### v1

# Beyond Classification: Towards Speech Emotion Reasoning with Multitask AudioLLMs

Abstract
Audio Large Language Models (AudioLLMs) have achieved strong results in semantic tasks like speech recognition and translation, but remain limited in modeling paralinguistic cues such as emotion. Existing approaches often treat emotion understanding as a classification problem, offering little insight into the underlying rationale behind predictions. In this work, we explore emotion reasoning, a strategy that leverages the generative capabilities of AudioLLMs to enhance emotion recognition by producing semantically aligned, evidence-grounded explanations. To support this in multitask AudioLLMs, we introduce a unified framework combining reasoning-augmented data supervision, dual-encoder architecture, and task-alternating training. This approach enables AudioLLMs to effectively learn different tasks while incorporating emotional reasoning. Experiments on IEMOCAP and MELD show that our approach not only improves emotion prediction accuracy but also enhances the coherence and evidential grounding of the generated responses.
[…]
tcb@breakable
Beyond Classification: Towards Speech Emotion Reasoning
with Multitask AudioLLMs
[…]
footnotetext: Contributed equally to this work; authors are listed in alphabetical order.
1 Introduction
Recent advancements in Audio Large Language Models (AudioLLMs) MERaLiON Team (2024); Tang et al. (2024); Chu et al. (2023); Hu et al. (2023); Das et al. (2024); D’efossez et al. (2024) have driven significant progress in spoken language understanding, particularly for tasks focused on semantic content such as automatic speech recognition (ASR), speech translation (ST), and spoken question answering (SQA). These models typically rely on large-scale audio-text alignment to align spoken inputs with textual outputs Ji et al. (2024); Held et al. (2024). However, current AudioLLMs are limited in modeling paralinguistic information, such as emotion, which is crucial for applications requiring emotionally aware or empathetic machine behavior Wang et al. (2024a); Sakshi et al. (2025); Ao et al. (2024).
Traditional emotion recognition approaches in speech primarily focus on categorical classification (e.g., predicting whether a speaker is angry or sad) Ma et al. (2024); Fu et al. (2025); Zhao et al. (2025). While effective for high-level emotion detection, such methods offer little interpretability or reasoning about why an emotion is being expressed. .
In this work, we leverage the generative capabilities of AudioLLMs to incorporate reasoning Ma et al. (2025); Xie et al. (2025) as a means to improve emotion recognition. Rather than treating emotion understanding as a purely discriminative task, we guide models to generate grounded, semantically aligned explanations that reflect both what is said (semantic content) and how it is said (paralinguistic cues). We categorize emotion recognition outputs into three types in Figure 1: (1) Label Only: direct classification (e.g. “The speaker is feeling angry”), with no explanation or grounding; (2) Interpretive Reasoning: explanation via paraphrased intent or inferred state (e.g. expressing frustration due to repeated failure); (3) Evidence-Grounded Reasoning: the most desirable form, which combines emotion labels with quoted utterances (e.g. “I’m not starting over again”) and interprets them to justify the emotional state.
Figure 1: Overview: Our proposed method leverages the generative capabilities of AudioLLM to go beyond classification, producing emotion labels alongside grounded, transcript-informed explanations that reflect the semantic and paralinguistic content of the input speech.
To this end, we propose a new multitask AudioLLM framework with multi-faceted strategies across data, architecture, and training. To guide the model’s generative capabilities, we construct reasoning-augmented supervision signals from transcript-aligned data, allowing the model to learn to produce emotion explanations grounded in both linguistic and paralinguistic evidence. Our model architecture employs a dual-encoder design that disentangles semantic and emotional representations. We also propose a task-alternating training strategy that separately optimizes the semantic and emotion encoders on their respective objectives, aiming to balance performance across tasks. Our framework is evaluated on benchmark datasets for emotion and sentiment recognition, namely IEMOCAP Busso et al. (2008) and MELD Poria et al. (2019), as well as ASR and SQA tasks. In summary, our main contributions are:
  • We propose a reasoning-augmented approach for speech emotion recognition, enabling AudioLLMs to generate semantically aligned, evidence-grounded explanations that enhance both interpretability and prediction accuracy.
  • We introduce a unified framework with multi-faceted strategies in data construction (reasoning target creation), architecture (dual-encoder design), and training (task-alternating training) for multitask AudioLLMs.
  • We conduct comprehensive experiments, showing that our approach effectively balances different task performances, improves emotion predictions with minimal effects on other tasks, and enables the coherence and grounding of generated emotion reasoning.
2 Related Works
2.1 AudioLLMs
Multimodal large language models (LLMs), including AudioLLMs MERaLiON Team (2024); Tang et al. (2024); Chu et al. (2023); Deshmukh et al. (2023); Hu et al. (2023); Das et al. (2024), commonly adopt a modular architecture comprising three core components: (1) a modality-specific encoder that extracts features from non-textual inputs, (2) a projection or adapter module that maps these features into a representation space compatible with the LLM’s tokenizer, and (3) a pretrained LLM that generates free-form text responses based on the projected modality tokens and natural language prompts. For instance, Qwen-Audio Chu et al. (2023) connects a Whisper-large-v2 Radford et al. (2023) speech encoder to the Qwen-7B Bai et al. (2023) language model. To capture richer audio representations, several models employ dual encoders that separately model semantic and acoustic information. SALMONN Tang et al. (2024) integrates Whisper-large-v2 and BEATs Chen et al. (2023) with Vicuna-13B Chiang et al. (2023), while WavLLM Hu et al. (2023) utilizes Whisper-large-v2 and WavLM-base Chen et al. (2022), interfaced with LLaMA-2-chat-7B Touvron et al. (2023).
Distillation approaches use LLMs to generate responses from speech transcriptions or metadata, such as gender and emotion, to train AudioLLMs. Kang et al. (2024) uses an LLM to generate responses to expressive speech prompts, Wang et al. (2024b) generate emotion-aware text continuations, and Lu et al. (2024a, b) generate detailed captions that reflect writing styles and tones.
Recent works explore enabling AudioLLMs to reason. Audio-CoT Ma et al. (2025) evaluates training-free chain-of-thought prompting, which requires AudioLLMs capable of general instruction following. Audio-Reasoner Xie et al. (2025) trains with a structured reasoning framework consisting of planning, captioning, reasoning and summarizing steps. Li et al. (2025) argues that the complex reasoning process in Audio-Reasoner may not be necessary, and the best practice remains an open research question.
Figure 2: Emotion reason extraction: We input the transcript, emotion label, and reasoning prompt into a speech-conditioned prompting template to elicit grounded and semantically aligned emotion explanations from the LLM. The Summarize prompt encourages interpretive reasoning based on the implied context, while the Elaborate prompt encourages evidence-grounded reasoning based on explicit cues from the transcript. For more coarse-grained sentiment reason extraction, replace the word "emotion" with "sentiment (positive, negative, neutral)" in the reasoning prompt.
2.2 Emotion recognition in AudioLLMs
Recent research in emotion recognition within AudioLLMs has explored a variety of strategies to enhance affective understanding from speech Bellver et al. (2024). These approaches leverage conversational context, paralinguistic cues, and ASR-generated transcripts to improve recognition accuracy. For instance, Sun et al. (2024) employs ASR and LLMs in a cascaded pipeline to transcribe and analyze emotional content, though such pipelines are susceptible to error propagation. SECap Xu et al. (2024) adopts contrastive and mutual information learning to disentangle semantic and emotional representations in speech. Fu et al. (2025) model speaker traits by prompting LLMs to infer emotional states based on listener responses. C²SER Zhao et al. (2025) combines Whisper and Emotion2Vec encoders with Chain-of-Thought prompting to inject contextual reasoning into emotion classification. SpeechCueLLM Wu et al. (2024) introduces descriptive cues, such as volume, pitch and speaking rate, into prompts to enrich LLM inputs with prosodic information.
In contrast to prior works, which focus primarily on improving emotion classification accuracy through architectural or input-level enhancements, our approach shifts the paradigm toward emotion reasoning. Rather than outputting a single emotion label, we leverage the generative capabilities of AudioLLMs to produce semantically coherent, evidence-grounded explanations for emotional states.
3 Proposed Method
Figure 3: Dual-encoder feature fusion and task-alternating training: We combine features from a general-purpose speech encoder and an emotion-centric encoder. Each encoder and its adapter are trained by alternating between speech-centric and emotion-centric tasks, enabling effective multitask learning with disentagled representations.
We propose a dual-encoder multitask AudioLLM framework that jointly models speech content and emotional reasoning. Our architecture integrates a general-purpose speech encoder and a specialized emotion-centric encoder, which are then connected to a large language model (LLM). To facilitate rich supervision, we introduce reasoning-augmented training targets derived from speech transcripts and emotion labels. Additionally, we adopt a task-alternating training strategy to ensure modular specialization and effective fusion of complementary features.
3.1 Emotion reason extraction
We introduce reasoning-augmented training targets that pair discrete emotion labels with natural language explanations. These explanations are derived through a prompting-based generation procedure, shown in Figure 2. Specifically, we construct a speech-conditioned prompting format that inputs the transcript, its associated emotion label, and a reasoning prompt into an LLM. We employ two distinct prompting strategies: the Summarize prompt encourages interpretive reasoning based on the broader implied context, while the Elaborate prompt guides the LLM to produce evidence-grounded justifications based on explicit cues from the transcript. The resulting explanations serve as supervision signals that teach the AudioLLM to associate emotional categories with meaningful linguistic and contextual cues, such that the resulting AudioLLM attains more interpretable and context-sensitive emotion understanding.
Using the generated targets, we construct question-answering training data by sampling questions from a curated pool designed to probe emotional understanding. These questions focus on the speaker’s affective state and examples include: "How would you interpret the speaker’s emotional state from their speech?", "What emotions do you think the speaker is expressing?", and "How would you describe the tone of the speaker’s voice?" We apply a similar approach when querying for coarser-grained sentiment, using broader questions that elicit the speaker’s overall positive, negative, or neutral disposition.
3.2 Dual-encoder feature fusion
The multitask AudioLLM framework we utilize consists of: (1) a dual-encoder architecture comprising a general-purpose speech encoder $E_{\text{speech}}$ and an emotion-centric encoder $E_{\text{emotion}}$, each designed to capture distinct aspects of the audio input; (2) a pair of lightweight adapter modules that project the encoder outputs into a shared latent space; and (3) a pretrained LLM that consumes the fused representation to generate free-form text outputs. The emotion-centric encoder serves to enhance emotion understanding and reasoning capabilities by introducing inductive biases specific to affective cues. An overview of the dual-encoder framework is shown in Figure 3.
We denote the dataset as $(\mathcal{A},\mathcal{T},\mathcal{Y})$, where $\mathcal{A}$ is the set of input audio signals, $\mathcal{T}$ is the set of corresponding text instructions, and $\mathcal{Y}$ is the set of output text responses. Given an audio input $a_{i}\in\mathcal{A}$ for the $i$-th training sample, we extract two types of audio embeddings: the speech encoder produces $z_{i}^{\text{speech}}=E_{\text{speech}}(a_{i})$, and the emotion-centric encoder produces $z_{i}^{\text{emotion}}=E_{\text{emotion}}(a_{i})$. We investigated different choices of emotion-centric encoder as described in Section 4 and fixed Whisper Radford et al. (2023), a widely used model for automatic speech recognition (ASR), as the speech encoder throughout our experiments. The utterances are zero-padded to 30 seconds, and the encoder embeddings have sequence length 1500. These encoder embeddings are passed through adapter modules to be reshaped and projected into a shared latent space. The speech encoder embeddings are transformed to have a sequence length of 100, while the emotion encoder embeddings are transformed to have a shorter sequence length of 10, emphasizing condensed emotion-specific representations as a complementary signal with minimal redundancy. We follow MERaLiON-AudioLLM MERaLiON Team (2024) in our implementation of the adapter modules: we concatenate the encoder embeddings across multiple time steps to reduce the sequence length, then pass the them through a multilayer perceptron (MLP) with two hidden layers and SiLU activation. The resulting audio token sequences are obtained as $\text{token}_{a_{i}}^{\text{speech}}$ and $\text{token}_{a_{i}}^{\text{emotion}}$. We concatenate these audio token sequences across the sequence dimension:
$\text{token}_{a_{i}}=\text{token}_{a_{i}}^{\text{speech}}\oplus_{s}\text{token}_{a_{i}}^{\text{emotion}}.$
Separately, we tokenize the text instruction $t_{i}\in\mathcal{T}$ as $\text{token}_{t_{i}}=\text{tokenizer}(t_{i})$. The audio and text tokens are then concatenated across the sequence dimension:
$\text{token}_{i}=\text{token}_{a_{i}}\oplus_{s}\text{token}_{t_{i}}.$
Finally, the concatenated tokens are fed into the LLM to generate the target response:
$\hat{y}_{i}=\text{LLM}(\text{token}_{i}).$
3.3 Task-alternating training
To ensure that each encoder specializes in its respective task, we employ a task-alternating training strategy, as illustrated in Figure 3. Specifically, the speech encoder and its adapter are trained on speech-centric tasks (e.g., spoken question answering, automatic speech recognition), while the emotion-centric encoder and its adapter are trained on emotion-centric tasks (e.g., emotion recognition with explanation generation). In each training round, only the encoder corresponding to the current task, its associated adapter, and the LLM LoRA parameters are updated, while the other encoder and adapter remain frozen. This alternating scheme enables disentangled yet complementary learning of speech and emotion representations. In the final epoch, we update all adapters and the LLM LoRA parameters across all tasks to enhance multimodal alignment.
4 Experiment Setup
4.1 Model implementation
We use Gemma-2-9B-IT Gemma Team et al. (2024) as the LLM for emotion reason extraction and in the AudioLLM framework. For each encoder, we utilize the encoder module from Whisper-Large-v3 Radford et al. (2023), which is a popular choice in existing AudioLLMs Tang et al. (2024); Chu et al. (2023); Hu et al. (2023); MERaLiON Team (2024). For the emotion-centric encoder, we also experiment with other options such as smaller-sized versions of Whisper, HuBERT Hsu et al. (2021) and Emotion2Vec Ma et al. (2024). For each adapter, we use a light-weight multilayer perceptron (MLP) with two hidden layers and SiLU activation function as in MERaLiON-AudioLLM MERaLiON Team (2024). We conduct multitask training with batch size 48 for 5 epochs on 8 H100 GPUs, using an AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and a learning rate of $5\times 10^{-5}$. The prompt template for LLM input takes the form:
“<audio_start> {audio tokens} <audio_end> {text instruction prompt}
4.2 Datasets
We conduct training and evaluation on two widely used benchmarks for emotion recognition (ER) and sentiment recognition (SR): IEMOCAP Busso et al. (2008) and MELD Poria et al. (2019). The IEMOCAP dataset comprises dyadic conversations between professional actors, where individual utterances are annotated with one of ten categorical emotion labels, namely anger, happiness, neutral, sadness, disgust, fear, surprise, frustration, excited and others. MELD is a multimodal dataset derived from the TV show Friends, containing audio, video, and text for multi-party conversations. In MELD-ER, each utterance is labeled with one of seven emotion classes, namely neutral, joy, disgust, sadness, surprise, anger and fear. We also include MELD-SR for sentiment recognition, where each utterance is labeled as positive, negative, or neutral, to evaluate the model’s ability to capture overall sentiment polarity in spoken contexts.
For semantic tasks, we utilize the Spoken Question Answering (SQA) tasks in MNSC Wang et al. (2025), a corpus centered on Singlish, a Creole language rooted in English. We select MNSC because the pre-trained encoders and LLM are unlikely to have been exposed to its linguistic patterns during training, reducing the risk of performance bias from prior exposure. For further analysis in Section 6, we experiment with additional SQA tasks, such as Spoken-SQuAD Lee et al. (2018) and SLUE-P2-SQA5 Shon et al. (2023), and automatic speech recognition (ASR) tasks such as MNSC ASR Wang et al. (2025) and LibriSpeech Panayotov et al. (2015).
Further data details and statistics are provided in the Appendix A.1.
4.3 Evaluation
We perform model evaluations on datasets in Section 4.2 using AudioBench Wang et al. (2024a) and follow its train-test splits to prevent data contamination. ASR tasks are evaluated with word error rate (WER), and remaining tasks are evaluated using LLM-as-a-Judge framework. Model outputs are assessed by Llama-3-70B-Instruct AI@Meta (2024) based on given scoring rubrics, and the scores are then normalized to 0-100 scale where higher scores reflect better performance.
For emotion and sentiment recognition, we grade each emotion prediction on a binary scale, where a score of 1 indicates semantic alignment with the ground-truth label. Since AudioLLMs generate open-ended responses, traditional metrics such as exact match may be insufficient. The LLM-as-a-Judge approach allows us to assess the factual correctness and relevance of model outputs in a more flexible and human-aligned manner.
We extend the LLM-as-a-Judge framework to evaluate the model’s evidence-grounded emotion reasoning quality. We extract direct quotes made in the model predictions and assess them using two key metrics: Groundedness Score and Relevance Score. Groundedness assesses how well the model’s extracted quotes align with the ground truth transcript, that is, whether they are directly quoted, faithfully paraphrased, or hallucinated. Relevance measures how effectively the extracted quotes support the annotated emotion label. For each prediction, we provide the Llama-3-70B-Instruct model judge with the ground-truth emotion label, speech transcript and extracted quotes, and instruct it to assign a score from 0 to 2 for each criterion based on a given structured scoring rubric. This evaluation captures both factual alignment and emotional interpretability of the model’s output. The evaluation prompt used is provided in the Appendix A.3.
5 Results
5.1 Effectiveness of reasoning-augmented training targets
We evaluate the impact of different supervision targets on emotion and sentiment recognition performance using a baseline AudioLLM equipped with a single Whisper-Large-v3 encoder. From results in Table 1, we observe that models trained with reasoning-augmented targets comprising both emotion labels and natural language explanations consistently outperform those trained with label-only supervision, with almost 20% improvement on average. This indicates that the inclusion of semantically rich and explanatory targets not only enhances interpretability but also leads to substantial gains in recognition capabilities, highlighting the effectiveness of grounding predictions in contextual reasoning. We provide examples of the different model predictions in Appendix A.2.
Training Targets IEMOCAP MELD-ER MELD-SR Avg Label Only (Original) 18.6 47.9 48.1 38.2 Interpretive Reasoning 60.8 52.6 60.1 57.8 Evidence-Grounded Reasoning 58.6 54.1 61.6 58.1
Table 1: Emotion and sentiment recognition performance of AudioLLM (with Whisper-Large-v3 encoder) trained on different supervision targets. Training with reasoning-augmented targets yields substantial performance improvements.
Figure 4: Strategies for feature fusion: We explore (a) fusion along the feature dimension, where features from both encoders are concatenated channel-wise, and (b) fusion along the sequence dimension, where features are concatenated token-wise across time steps.
[…]
Table 2: Performance comparison of different methods for dual-encoder feature fusion and multitask training. Concat Dim "None" indicates the single-encoder baseline. "Joint" indicates that all tasks are trained together on all sets of encoder + adapter for 5 epochs. "Alt $x$ epoch(s)" refers to alternating training of speech-centric and emotion-centric tasks on their respective encoder + adapter every $x$ epoch(s), up to 4 epochs of data, then training all tasks on all adapters at the final epoch.
[…]
Table 3: Performance comparison of different choices of emotion-centric encoder in the dual-encoder architecture. "Alt $x$ epoch(s)" refers to alternating training of speech-centric and emotion-centric tasks on their respective encoder + adapter every $x$ epoch(s), up to 4 epochs of data, then training all tasks on all adapters at the final epoch.
5.2 Effectiveness of dual-encoder feature fusion and training
Besides IEMOCAP and MELD, we train and evaluate on MNSC SQA Part 3-6 to assess the effects of our proposed method on non-emotion-centric tasks. Table 2 presents a systematic comparison of key design choices, including model architecture, feature fusion strategies, and task-alternating training. We use Whisper-Large-v3 for both encoders in the dual-encoder architectures. We observe that
  1. 1.
    Dual-encoder architectures can outperform single-encoder baselines, suggesting that incorporating complementary representations enhances overall performance;
  2. 2.
    Concatenation along the sequence dimension yields slightly better results than concatenation along the feature dimension, likely due to better preservation of temporal structure. In the latter, both speech and emotion embeddings are reshaped to length 100, concatenated along the feature dimension, and then passed through a single adapter module. The two types of concatenation are illustrated in Figure 4;
  3. 3.
    Task-alternating training leads to improved performance, particularly for emotion and sentiment recognition, compared to joint multitask training.
In Table 3, we further compare multitask performance across different choices of emotion-centric encoders. We adopt a dual-encoder architecture with features concatenated along the sequence dimension, and apply task-alternating training. Our observations include:
  1. 1.
    Each round of task-specific training must be sufficiently long to ensure model convergence;
  2. 2.
    Encoder selection has a notable impact on performance. Specifically, Emotion2Vec+ Large, which is pre-trained for emotion recognition, provides more relevant features for emotion-centric tasks, leading to improved emotion-centric and overall performance.
5.3 Quality of emotion reasoning responses
Table 4 presents the evaluation of evidence-grounded reasoning in the responses generated by our AudioLLM. Across all evaluated datasets, over 49% of the model’s responses explicitly include direct quotes from the speech content, serving as supporting evidence for reasoning or interpretation. The model achieves high groundedness scores, averaging 82.8%, indicating that most quotes are faithful to the original speech transcript. Relevance scores average 65.3%, suggesting that the majority of quotes meaningfully support the ground-truth emotion label, though there remains room for improvement, particularly on the MELD dataset.
Scores IEMOCAP MELD-ER MELD-SR Avg Quotation 73.8 49.2 49.6 57.5 Groundedness 90.6 79.2 78.5 82.8 Relevance 71.9 60.0 64.1 65.3
Table 4: Evaluation of model predictions with evidence-grounded reasoning predictions. Quotation Score measures the percentage of predictions containing at least one extractable quote. Groundedness and Relevance scores (0–100 scale) assess alignment with the transcript and support for the ground-truth emotion label, respectively.
5.4 Comparison with other models
We compare the emotion and sentiment recognition performance with end-to-end AudioLLMs evaluated in AudioBench: WavLLM Hu et al. (2023), Qwen2-Audio-7B-Instruct Chu et al. (2024), Phi-4-Multimodal-Instruct Abouelenin et al. (2025), MERaLiON-AudioLLM MERaLiON Team (2024), Qwen-Audio-Chat Chu et al. (2023), and SALMONN Tang et al. (2024). We also compare with cascaded models that process speech in sequential stages by converting audio to text using an automatic speech recognition module before feeding the transcript into a large language model: Whisper-Large-v2 with SEA-LIONv3 Ng et al. (2025), and Whisper-Large-v3 with Llama-3-8B-Instruct AI@Meta (2024). From Table 5, our proposed AudioLLM trained with reasoning-augmented targets achieves the best performance across IEMOCAP, MELD-ER and MELD-SR.
[…]
Table 5: Emotion and sentiment recognition performance of end-to-end AudioLLMs and cascaded models.
We compare the performance of AudioLLM with emotion classifier Emotion2Vec+ Large in Table 6 on MELD. AudioLLM trained with reasoning-augmented targets achieves the highest weighted average accuracy (53.0%), outperforming both the label-only variant (47.1%) and Emotion2Vec+ Large (44.7%). However, the unweighted average remains comparable across models, suggesting that rare classes, such as disgust and fear, remain challenging.
It is important to note that comparisons with existing models are not controlled for differences in training data, which may influence performance outcomes.
Class Num Samples Emotion2Vec+ large AudioLLM - Label Only AudioLLM - Reasoning Neutral 1256 54.2 64.7 83.8 Joy 402 54.5 41.5 36.3 Disgust 68 0.0 8.8 13.2 Sadness 208 32.2 23.1 15.9 Surprise 281 38.8 28.8 26.0 Anger 345 26.1 32.2 31.3 Fear 50 2.0 6.0 8.0 Unwt Avg 2610 29.7 29.3 29.2 Wt Avg 2610 44.7 47.1 53.0
Table 6: Class-wise emotion recognition performance comparison of Emotion2Vec+ Large classifier vs. AudioLLM (with Emotion2Vec+ Large emotion-centric encoder) trained on Label Only or Reasoning targets on MELD-ER. The unweighted average (Unwt Avg) treats all classes equally regardless of sample size, while the weighted average (Wt Avg) reflects the sample distribution by weighting each class accordingly.
6 Further Analysis
We further investigate whether emotion understanding capabilities can be introduced into a model that was not originally trained for these tasks. Starting from a base AudioLLM without any emotion-specific supervision, we explore adding an emotion-centric encoder to the architecture. We train the emotion-centric encoder Emotion2Vec+ Large, adapter and LLM LoRA on emotion-centric tasks, then finetune the adapters and LLM LoRA on all tasks.
Table 7 presents the effect of incorporating emotion supervision into a base AudioLLM trained on different upstream tasks. Across both training configurations, we observe substantial gains in emotion and sentiment recognition, with improvements of +12.3 and +22.4 points, respectively. This enhancement in emotional understanding comes with a slight degradation in the model’s performance on the original tasks. For instance, SQA performance remains comparable, and ASR performance is largely preserved or slightly improved. These results highlight that emotional capabilities can be effectively injected into a multimodal model without sacrificing its existing competencies.
Model ER/SR ($\uparrow$) SQA ($\uparrow$) ASR ($\downarrow$) Base AudioLLM 44.1 56.0 19.5 + Emotion Supervision 56.4 54.1 19.6
(a) Base AudioLLM trained on MNSC SQA Part 3-6 and MNSC ASR Part 3-6.
Model ER/SR ($\uparrow$) SQA ($\uparrow$) ASR ($\downarrow$) Base AudioLLM 35.6 80.3 3.8 + Emotion Supervision 58.0 79.0 3.6
(b) Base AudioLLM trained on Spoken-SQuAD and SLUE for SQA, and LibriSpeech Clean and Other splits for ASR.
Table 7: Effect of adding emotion supervision on a trained base AudioLLM. Adding emotion supervision improves emotion understanding with slight compromise on the performance of other tasks
7 Conclusion
In this work, we propose a unified framework that brings emotion reasoning into multitask AudioLLMs, combining dual encoders, reasoning-augmented supervision, and task-alternating training. Our method improves emotion and sentiment recognition and enables the generation of evidence-grounded explanations, as demonstrated on IEMOCAP and MELD benchmark datasts. This work highlights the potential of generative AudioLLMs for more interpretable and emotionally aware speech understanding.
8 Limitations
While our framework significantly improves both emotion recognition and explanation capabilities in AudioLLMs, several limitations remain. The explanation generation quality can vary across emotions and speaker styles, especially for subtle or ambiguous affective states. Moreover, the quality of extracted emotion reasoning is dependent on the capabilities of the teacher LLM, which may introduce biases or inaccuracies in supervision. The current benchmarks for emotional reasoning are limited in diversity and scale, highlighting the need for more comprehensive evaluation datasets that capture a wider range of emotional expressions and contextual richness.
9 Ethics Statement
All datasets used in our study, including IEMOCAP and MELD, are publicly available and widely used in the research community. However, we emphasize that caution is necessary when deploying emotion-aware AI systems in real-world or sensitive contexts, as misinterpretation of emotional cues may lead to unintended consequences. Ensuring transparency, user consent, and appropriate safeguards is critical when applying these technologies beyond academic settings.
Acknowledgements
This research is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore.
Appendix A Experiment Details
A.1 Datasets
We conduct training and evaluation on two widely used benchmarks for emotion recognition (ER) and sentiment recognition (SR): IEMOCAP Busso et al. (2008) and MELD Poria et al. (2019). IEMOCAP is made available under a custom non-commercial research license, and MELD is distributed under the GNU General Public License v3.0 (GPL-3.0). Since IEMOCAP lacks a predefined train-test split, we adopt the 90-10 split defined in AudioBench Wang et al. (2024a), with 9035 training samples (anger: 1140, disgust: 2, excited: 1816, fear: 98, frustration: 2608, happiness: 588, neutral: 1539, other: 23, sad: 1120, surprise: 101) and 1004 test samples (anger: 129, disgust: 0, excited: 160, fear: 9, frustration: 309, happiness: 68, neutral: 187, other: 3, sad: 130, surprise: 9). MELD has 9988 training samples (anger: 1109, disgust: 271, fear: 268, joy: 1743, neutral: 4709, sad: 683, surprise: 1205) and 2610 test samples (anger: 345, disgust: 68, fear: 50, joy: 402, neutral: 1256, sadness: 208, surprise: 281).
For semantic tasks, we utilize the Multitask National Speech Corpus (MNSC) Wang et al. (2025), specifically SQA Part 3-6 and ASR Part 3-6, released under the Singapore Open Data License. We also use Spoken-SQuAD Lee et al. (2018) released under CC-BY-SA-4.0 License, SLUE-P2-SQA5 Shon et al. (2023) which is a collection of datasets released under CC-BY-SA-4.0 License and Apache License 2.0, and LibriSpeech Panayotov et al. (2015) released under CC-BY-4.0 License.
All experiments in this work respect the respective licenses and usage terms of the datasets.
A.2 AudioLLM responses
We conduct a qualitative analysis of model predictions across different reasoning formats on the IEMOCAP, MELD-ER and MELD-SR datasets. Each example includes the transcript, ground-truth label, and model-generated outputs under three supervision types: label-only, interpretive reasoning, and evidence-grounded reasoning. As shown in Table 8, 9 and 10, label-only responses often fail to capture the correct emotional nuance, defaulting to neutral predictions even when the emotion is apparent. In contrast, interpretive and evidence-grounded reasoning better align with the ground truth, offering richer justifications and improved emotion recognition. Notably, evidence-grounded reasoning demonstrates superior clarity by explicitly linking speech content and affective cues to the predicted emotion.
Example 1
Transcript: Is he an actor too, or-?
Ground Truth Emotion Label: Neutral
[…]
Interpretive Reasoning: The speaker is asking if someone was an actor too, implying a shared experience or connection. They seem neutral in their tone, neither excited nor disappointed.
Evidence-Grounded Reasoning: The speaker’s tone is neutral, suggesting they are neither expressing strong positive nor negative emotions. They are simply seeking clarification or information. The phrase "Is he an actor too?" implies a previous context where someone was identified as an actor. The speaker is now inquiring if another person also shares this profession.
Example 2
Transcript: Oh, I will never forget his face. Ooh-rah [LAUGHTER].
Ground Truth Emotion Label: Excited
[…]
Interpretive Reasoning: The speaker is laughing and reminiscing about someone’s face, implying a positive and humorous memory. They express excitement and amusement.
Evidence-Grounded Reasoning: The speaker is clearly excited and perhaps a little bit nostalgic. Here’s why: "Oh, I will never forget his face": This phrase suggests a strong memory, likely positive given the excited tone. Laughter: Laughter often accompanies happy memories and a sense of lightheartedness. The speaker is reminiscing about someone and a specific moment, and the excitement suggests that the memory is a pleasant one.
Example 3
Transcript: You’ve got to tell me.
Ground Truth Emotion Label: Frustration
[…]
Interpretive Reasoning: The speaker is frustrated and wants the listener to tell them something.
Evidence-Grounded Reasoning: The speaker is frustrated. They are likely feeling impatient, annoyed, or even angry because they feel like something is not being done or explained properly. The phrase "You’ve got to tell me" suggests a sense of urgency and a demand for information.
Table 8: IEMOCAP: Qualitative comparison of emotion predictions. The first three samples in the evaluation dataset are presented. Each example includes the transcript, ground-truth emotion label, and model predictions with their inferred emotions. Emotion labels are in blue, and quotes from the speech content are in orange. Interpretive and evidence-grounded reasoning better align with the ground truth, offering richer justifications and improved emotion recognition.
Example 1
Transcript: Yeah, I’m not in that.
Ground Truth Emotion Label: Neutral
[…]
Interpretive Reasoning: The speaker is denying being in something, and they sound neutral about it.
Evidence-Grounded Reasoning: The speaker’s tone is neutral, indicating they are not experiencing any strong emotions. They are simply stating a fact without any particular emphasis or feeling behind their words.
Example 2
Transcript: He’s a dropper!
Ground Truth Emotion Label: Anger
[…]
Interpretive Reasoning: The speaker is expressing anger and frustration towards someone they perceive as a "dropper," likely someone who is unreliable or lets them down.
Evidence-Grounded Reasoning: The speaker’s tone is neutral, but the statement itself carries a strong implication. While the speaker doesn’t express anger or disgust directly, the phrase "He’s a dropper!" suggests a negative judgment about the person being referred to. It implies that this person is unreliable, perhaps someone who makes promises they don’t keep or fails to deliver on expectations.
Example 3
Transcript: Yes!
Ground Truth Emotion Label: Joy
[…]
Interpretive Reasoning: The speaker said "Yes." The emotion conveyed is neutral.
Evidence-Grounded Reasoning: The speaker’s tone is neutral, indicating they are not experiencing any strong positive or negative emotions. They are likely calm, composed, and objective in their state of mind.
Table 9: MELD-ER: Qualitative comparison of emotion predictions. The first three samples in the evaluation dataset are presented. Each example includes the transcript, ground-truth emotion label, and model predictions with their inferred emotions. Emotion labels are in blue, and quotes from the speech content are in orange. Interpretive and evidence-grounded reasoning provide richer interpretation of speaker emotion by additionally making use of the speech content.
Example 1
Transcript: Yeah, I’m not in that.
Ground Truth Emotion Label: Neutral
[…]
Interpretive Reasoning: The speaker is stating that they are not in a particular location or situation. They express this in a neutral tone, indicating no strong feelings about the matter.
Evidence-Grounded Reasoning: The speaker’s tone is neutral, indicating they are not expressing strong positive or negative emotions. They are simply stating a fact without any noticeable inflection or emphasis.
Example 2
Transcript: He’s a dropper!
Ground Truth Emotion Label: Negative
[…]
Interpretive Reasoning: The speaker is expressing a negative sentiment towards someone they perceive as a "dropper," which likely implies a negative trait or action.
Evidence-Grounded Reasoning: The sentiment is negative. The speaker’s tone suggests disapproval or disgust towards the person being referred to as a "dropper." The word "dropper" itself often carries a negative connotation, implying someone who is unreliable, messy, or in some way undesirable.
Example 3
Transcript: Yes!
Ground Truth Emotion Label: Positive
[…]
Interpretive Reasoning: The speaker said "Yes." The sentiment is neutral.
Evidence-Grounded Reasoning: The sentiment is positive. The speaker’s enthusiastic "Yes!" conveys a strong sense of agreement, excitement, or approval.
Table 10: MELD-SR: Qualitative comparison of sentiment predictions. The first three samples in the evaluation dataset are presented. Each example includes the transcript, ground-truth sentiment label, and model predictions with their inferred sentiments. Sentiment labels are in blue, and quotes from the speech content are in orange. Interpretive and evidence-grounded reasoning better align with the ground truth, offering richer justifications and improved sentiment recognition.
A.3 Evaluation
We follow AudioBench’s LLM-as-a-Judge framework to evaluate the model’s evidence-grounded emotion reasoning quality. We extract direct quotes made in the model predictions and assess them using two key metrics: Groundedness Score and Relevance Score. For each prediction, we provide the Llama-3-70B-Instruct model judge with the ground-truth emotion label, speech transcript and extracted quotes, and instruct it to assign a score from 0 to 2 for each criterion based on a given structured scoring rubric. The scores are then normalized to 0-100 scale. The evaluation prompt used is shown in Figure 5.
Evaluation Prompt Template
Figure 5: Evaluation prompt used for assessing groundedness and relevance of extracted emotional evidence.